OpenStudy (anonymous):

simplify 1/cos^2(x)-(sin^4(x)/cos^4(x))

OpenStudy (zehanz):

Try using sec(x)=1/cos(x) and tan(x)=sin(x)/cos(x).

OpenStudy (anonymous):

can it be done with just cos and sin?

OpenStudy (zehanz):

It can, but I can't make it much simpler with these:\[\frac{ 1 }{ \cos^2x }-\frac{ \sin^4x }{ \cos^4x }=\frac{ \cos^2x }{ \cos^4x }-\frac{ \sin^4x }{ \cos^4x }=\frac{ \cos^2x-\sin^4x }{ \cos^4x }\]From here on I get only more instead of less ;)

OpenStudy (anonymous):

ok, thanks.

OpenStudy (zehanz):

YW! The hint I gave in the beginning leads to:\[\sec^2x-\tan^4x\]Looks simpler...